# Finite-Sample Analysis of Policy Evaluation for Robust Average Reward Reinforcement Learning

### v1

@@ A.2 Proof of Theorem 2
$\leq\sum_{a\in\mathcal{A}}\pi(a|s)\max_{p\in\mathcal{P}^{a}_{s}}\Bigg{[}\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,a)V_{1}(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,a)V_{2}(s^{\prime})\Bigg{]}$ (44)
Where the last inequality comes from the fact that $\min f(x)-\min g(x)\leq\max(f(x)-g(x))$. denote $\tilde{p}(V_{1},V_{2})=\operatornamewithlimits{argmax}_{p\in\mathcal{P}^{a}_{s}}[\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,a)V_{1}(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,a)V_{2}(s^{\prime})]$, we further obtain
$\mathbf{T}_{g}(V_{1})(s)-\mathbf{T}_{g}(V_{2})(s)$ $\leq\sum_{a\in\mathcal{A}}\pi(a|s)\Bigg{[}\sum_{s^{\prime}\in\mathcal{S}}\tilde{p}(s^{\prime}|s,a)V_{1}(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}}\tilde{p}(s^{\prime}|s,a)V_{2}(s^{\prime})\Bigg{]}$
$\leq\sum_{a\in\mathcal{A}}\pi(a|s)\sum_{s^{\prime}\in\mathcal{S}}\tilde{p}(s^{\prime}|s,a)[V_{1}(s^{\prime})-V_{2}(s^{\prime})]$ (45)
Since $\tilde{p}(V_{1},V_{2})\in\mathcal{P}$ for all $V_{1},V_{2}$, the above problem transforms to the same as the non-robust setting. Thus, by Lemma 1, there exists $\beta(V_{1},V_{2})\in(0,1)$ such that
${\|\mathbf{T}_{g}(V_{1})-\mathbf{T}_{g}(V_{2})\|}_{\mathrm{sp}}\leq\left\lVert\sum_{a\in\mathcal{A}}\pi(a|s)\sum_{s^{\prime}\in\mathcal{S}}\tilde{p}(s^{\prime}|s,a)[V_{1}(s^{\prime})-V_{2}(s^{\prime})]\right\rVert_{\mathrm{sp}}\leq\beta(V_{1},V_{2}){\|V_{1}-V_{2}\|}_{\mathrm{sp}}$ (46)
Give that $\mathcal{P}$ is compact, the supremum is attainable and there exist $\gamma\in(0,1)$ such that for all $V_{1},V_{2}$, we have ${\|\mathbf{T}_{g}(V_{1})-\mathbf{T}_{g}(V_{2})\|}_{\mathrm{sp}}\leq\gamma{\|V_{1}-V_{2}\|}_{\mathrm{sp}}$, which concludes the proof.
Appendix B Biased Semi-Norm Stochastic Approximation
@@
We perform analysis of the biased‐noise extension to the span semi-norm stochastic approximation (SA) problem by constructing a smooth convex semi-Lyapunov function for forming the negative drift (Zhang et al., 2021) and using properties in dual norms for managing the bias.
B.1 Setup and Notation.
@@
In this section, we override the notation of the span semi-norm by re-writing it as the sup-norm to the equivalence class of constant vectors. For any norm $\|\cdot\|_{c}$ and equivalent class $\delta_{\overline{E}}$, define the indicator function $\delta_{\overline{E}}$ as
@@
then by (Zhang et al., 2021), the semi-norm induced by norm $\|\cdot\|_{c}$ and equivalent class $\overline{E}$ is the infimal convolution of $\|\cdot\|_{c}$ and the indicator function $\delta_{\overline{E}}$ can be defined as follows
@@
Where $\ast_{\inf}$ denotes the infimal convolution operator. Throughout the remaining section, we let $\overline{E}\coloneqq\{c\mathbf{e}:c\in\mathbb{R}\}$ with $\mathbf{e}$ being the all-ones vector. Then from (Gupta et al., 2015), we have for any $x$,
$\|x\|_{\mathrm{sp}}=2\|x\|_{\infty,\overline{E}}.$ (49)
@@
$\|H(x)-H(y)\|_{\infty,\overline{E}}\leq\gamma\|x-y\|_{\infty,\overline{E}},\quad\gamma\in(0,1),\quad\forall x,y$ (50)
@@
  • $\mathbb{E}[\,\|w^{t}\|_{\infty,\overline{E}}^{2}|\mathcal{F}^{t}]\leq A+B\|x^{t}-x^{*}\|_{\infty,\overline{E}}^{2}$ (In the robust average reward TD case, $B=0$).
  • $\bigl{\|}\mathbb{E}[w^{t}|\mathcal{F}^{t}]\bigr{\|}_{\infty,\overline{E}}\leq\varepsilon_{\mathrm{bias}}$.
@@
Note that beside the bias in the noise, the above formulation and assumptions are identical to the unbiased setups in Section B of (Zhang et al., 2021). Thus, we emphasize mostly on managing the bias.
B.2 Semi‐Lyapunov $M_{\overline{E}}(\cdot)$ and Smoothness.
By (Zhang et al., 2021, Proposition 1–2), using the Moreau envelope function $M(x)$ in Definition 2.2 of (Chen et al., 2020), we define
@@
$c_{l}M_{\overline{E}}(x)\leq\frac{1}{2}\|x\|_{\infty,\overline{E}}^{2}\leq c_{u}M_{\overline{E}}(x),$ (52)

### v2

@@ C.2 Proof of Theorem 2
$\leq\sum_{a\in\mathcal{A}}\pi(a|s)\max_{p\in\mathcal{P}^{a}_{s}}\Bigg{[}\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,a)V_{1}(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,a)V_{2}(s^{\prime})\Bigg{]}$
$\leq\sum_{a}\pi(a|s)\sum_{s^{\prime}}\tilde{p}_{(V_{1},V_{2})}(s^{\prime}|s,a)\bigl{[}V_{1}(s^{\prime})-V_{2}(s^{\prime})\bigr{]}$ (40)
where $\tilde{p}_{(V_{1},V_{2})}=\arg\max_{p\in\mathcal{P}^{a}_{s}}[\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,a)V_{1}(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,a)V_{2}(s^{\prime})]$ and each $\tilde{p}_{(V_{1},V_{2})}\in\mathcal{P}$ for all $V_{1},V_{2}$. We now discuss the construction of the desired semi-norm $\|\cdot\|_{\mathcal{P}}$.
For any $\mathsf{P}\in\mathcal{P}$, denote $\mathsf{P}^{\pi}$ as the transition matrix under policy $\pi$ and the unique stationary distribution $d^{\pi}_{\mathsf{P}}$, and denote $E_{\mathsf{P}}$ as the matrix with all rows being identical to $d^{\pi}_{\mathsf{P}}$. We further define $Q_{\mathsf{P}}^{\pi}=\mathsf{P}^{\pi}-E_{\mathsf{P}}$. Thus, by Lemma 4 we have $\rho(Q_{\mathsf{P}}^{\pi})<1$ for all $\mathsf{P}\in\mathcal{P}$. Since $\mathcal{P}$ is compact, denote $r^{*}$ as the supremum of the spectral radius of all $Q_{\mathsf{P}}^{\pi}$, we have:
$r^{*}\coloneqq\sup_{\mathsf{P}\in\mathcal{P}}\rho(Q_{\mathsf{P}}^{\pi})<1$ (41)
We follow similar process for constructing our desired semi-norm $\|\cdot\|_{\mathcal{P}}$ as in Appendix C.1 by first constructing a norm such that all $Q_{\mathsf{P}}^{\pi}$ are strictly less then one under that norm. We choose $\alpha\in(r^{*},1)$ and we follow the approach in Wirth (2002) by constructing an extremal norm $\|\cdot\|_{\rm ext}$ as follows:
$\|x\|_{\rm ext}\coloneqq\sup_{k\geq 0}\sup_{Q_{1},\ldots,Q_{k}\in Q^{\pi}_{\mathcal{P}}}\alpha^{-k}\|Q_{k}Q_{k-1}\ldots Q_{1}x\|_{2}\quad\text{where}\quad Q^{\pi}_{\mathcal{P}}=\{Q^{\pi}_{\mathsf{P}}:\mathsf{P}\in\mathcal{P}\}$ (42)
Note that we follow the convention that $\|Q_{k}Q_{k-1}\ldots Q_{1}x\|_{2}=\|x\|_{2}$ when $k=0$.
Lemma 6.
The operator $\|\cdot\|_{\rm ext}$ is a valid norm with $\|Q_{\mathsf{P}}^{\pi}\|_{\rm ext}<1$ for all $\mathsf{P}\in\mathcal{P}$
Proof.
We first prove that $\|\cdot\|_{\rm ext}$ is bounded. Following Lemma 10 and choosing $\lambda\in(r^{*},\alpha)$, then there exist a positive constant $C<\infty$ such that
$\|Q_{k}Q_{k-1}\ldots Q_{1}\|_{2}\leq C\lambda^{k}$ (43)
Hence for each $k$ and for all $x\in\mathbb{R}^{S}$,
$\alpha^{-k}\|Q_{k}Q_{k-1}\ldots Q_{1}x\|_{2}\leq\alpha^{-k}C\lambda^{k}\|x\|_{2}=C\left(\frac{\lambda}{\alpha}\right)^{-k}\|x\|_{0}\longrightarrow 0\quad\text{as}\quad k\rightarrow\infty$ (44)
Thus the double supremum in (42) is over a bounded and vanishing sequence, so $\|\cdot\|_{\rm ext}$ bounded.
To check that $\|\cdot\|_{\rm ext}$ is a valid norm, note that if $x=\bf{0}$, $\|x\|_{\rm ext}$ is directly 00. On the other hand, if $\|x\|_{\rm ext}=0$, we have
$\sup_{k=0}\sup_{Q_{1},\ldots,Q_{k}\in Q^{\pi}_{\mathcal{P}}}\alpha^{-k}\|Q_{k}Q_{k-1}\ldots Q_{1}x\|_{2}=\|x\|_{2}=0\Rightarrow x=\bf{0}$ (45)
Regarding homogeneity, observe that for any $c\in\mathbb{R}$ and $x\in\mathbb{R}^{S}$,
$\|cx\|_{\rm ext}=\sup_{k\geq 0}\sup_{Q_{1},\ldots,Q_{k}\in Q^{\pi}_{\mathcal{P}}}\alpha^{-k}\|Q_{k}Q_{k-1}\ldots Q_{1}(cx)\|_{2}=|c|\|x\|_{\rm ext}$ (46)
Regarding triangle inequality, using $\|Q_{k}\ldots Q_{1}(x+y)\|_{2}\leq\|Q_{k}\ldots Q_{1}x\|_{2}+\|Q_{k}\ldots Q_{1}y\|_{2}$ for any $x,y\in\mathbb{R}^{S}$, we obtain,
$\|x+y\|_{\rm est}=\sup_{k\geq 0}\sup_{Q_{1},\ldots,Q_{k}\in Q^{\pi}_{\mathcal{P}}}\alpha^{-k}\|Q_{k}Q_{k-1}\ldots Q_{1}(x+y)\|_{2}\leq\|x\|_{\rm est}+\|y\|_{\rm est}$ (47)
For any $\mathsf{P}\in\mathcal{P}$, we have
$\|Q_{\mathsf{P}}^{\pi}x\|_{\rm ext}$ $=\sup_{k\geq 0}\sup_{Q_{1},\ldots,Q_{k}\in Q^{\pi}_{\mathcal{P}}}\alpha^{-k}\|Q_{k}Q_{k-1}\ldots Q_{1}(Q_{\mathsf{P}}^{\pi}x)\|_{2}$
$\leq\sup_{k\geq 1}\sup_{Q_{1},\ldots,Q_{k}\in Q^{\pi}_{\mathcal{P}}}\alpha^{-(k-1)}\|Q_{k}Q_{k-1}\ldots Q_{1}x\|_{2}$
$=\alpha\sup_{k\geq 1}\sup_{Q_{1},\ldots,Q_{k}\in Q^{\pi}_{\mathcal{P}}}\alpha^{-k}\|Q_{k}Q_{k-1}\ldots Q_{1}x\|_{2}$
$\leq\alpha\sup_{k\geq 0}\sup_{Q_{1},\ldots,Q_{k}\in Q^{\pi}_{\mathcal{P}}}\alpha^{-k}\|Q_{k}Q_{k-1}\ldots Q_{1}x\|_{2}$
$=\alpha\|x\|_{\rm ext}$ (48)
Since $\mathsf{P}$ is arbitrary, (6) implies that for any $\mathsf{P}\in\mathcal{P}$,
$\|Q_{\mathsf{P}}^{\pi}\|_{\rm ext}=\sup_{x\neq\bf{0}}\frac{\|Q_{\mathsf{P}}^{\pi}x\|_{\rm ext}}{\|x\|_{\rm ext}}\leq\alpha<1$ (49)
We now follow the same method as (36) to construct the semi-norm $\|\cdot\|_{\mathcal{P}}$. Define the operator $\|\cdot\|_{\mathcal{P}}$ as
$\|x\|_{\mathcal{P}}\coloneqq\sup_{\mathsf{P}\in\mathcal{P}}\bigl{\|}Q_{\mathsf{P}}^{\pi}x\bigr{\|}_{\rm ext}+\epsilon\inf_{c\in\mathbb{R}}\bigl{\|}x-c\mathbf{e}\bigr{\|}_{\rm ext}$ (50)
where $0<\epsilon<1-\alpha$.
Lemma 7.
The operator $\|\cdot\|_{\mathcal{P}}$ is a valid semi-norm with kernel being exactly $\{c\mathbf{e}:c\in\mathbb{R}\}$. Furthermore, for all $x\in\mathbb{R}^{S}$, we have $\bigl{\|}\mathsf{P}^{\pi}x\bigr{\|}_{\mathcal{P}}\leq(\alpha+\epsilon)\bigl{\|}x\bigr{\|}_{\mathcal{P}}$.
Proof.
Regarding positive homogeneity and nonnegativity, for any scalar $\lambda$ and $x\in\mathbb{R}^{S}$,
$\|\lambda x\|_{\mathcal{P}}=\sup_{\mathsf{P}\in\mathcal{P}}\bigl{\|}Q_{\mathsf{P}}^{\pi}(\lambda x)\bigr{\|}_{\rm ext}+\epsilon\inf_{c\in\mathbb{R}}\bigl{\|}\lambda x-c\mathbf{e}\bigr{\|}_{\rm ext}=|\lambda|\sup_{\mathsf{P}\in\mathcal{P}}\bigl{\|}Q_{\mathsf{P}}^{\pi}x\bigr{\|}_{\rm ext}+\epsilon|\lambda|\inf_{c\in\mathbb{R}}\bigl{\|}x-c\mathbf{e}\bigr{\|}_{\rm ext}=|\lambda|\|x\|_{\rm ext}$
and $\|x\|_{\rm ext}\geq 0$. Regarding triangle inequality, for any $x,y\in\mathbb{R}^{S}$, note that for any $\mathsf{P}\in\mathcal{P}$,
$\|Q^{\pi}_{\mathsf{P}}(x+y)\|_{\rm ext}\leq\|Q^{\pi}_{\mathsf{P}}x\|_{\rm ext}+\|Q^{\pi}_{\mathsf{P}}y\|_{\rm ext}$ (51)
Taking supremum over $\mathsf{P}$ on both sides yields
$\sup_{\mathsf{P}\in\mathcal{P}}\|Q^{\pi}_{\mathsf{P}}(x+y)\|_{\rm ext}\leq\sup_{\mathsf{P}\in\mathcal{P}}\|Q^{\pi}_{\mathsf{P}}x\|_{\rm ext}+\sup_{\mathsf{P}\in\mathcal{P}}\|Q^{\pi}_{\mathsf{P}}y\|_{\rm ext}$ (52)
Thus, we have
$\|x+y\|_{\mathcal{P}}$ $=\sup_{\mathsf{P}\in\mathcal{P}}\bigl{\|}Q_{\mathsf{P}}^{\pi}(x+y)\bigr{\|}_{\rm ext}+\epsilon\inf_{c}\bigl{\|}x+y-c\mathbf{e}\bigr{\|}_{\rm ext}$
$\leq\sup_{\mathsf{P}\in\mathcal{P}}\|Q^{\pi}_{\mathsf{P}}x\|_{\rm ext}+\sup_{\mathsf{P}\in\mathcal{P}}\|Q^{\pi}_{\mathsf{P}}y\|_{\rm ext}+\epsilon\inf_{a,b}\bigl{\|}x-a\mathbf{e}+y-b\mathbf{e}\bigr{\|}_{\rm ext}$
$\leq\sup_{\mathsf{P}\in\mathcal{P}}\|Q^{\pi}_{\mathsf{P}}x\|_{\rm ext}+\sup_{\mathsf{P}\in\mathcal{P}}\|Q^{\pi}_{\mathsf{P}}y\|_{\rm ext}+\epsilon\inf_{a}\bigl{\|}x-a\mathbf{e}\bigr{\|}_{\rm ext}+\epsilon\inf_{b}\bigl{\|}y-b\mathbf{e}\bigr{\|}_{\rm ext}$
$=\|x\|_{\mathcal{P}}+\|y\|_{\mathcal{P}}.$
Regarding the kernel, if $x=k\mathbf{e}$ for some $k\in\mathbb{R}$, then similar to (5), we have
$\|x\|_{\mathcal{P}}=\sup_{\mathsf{P}\in\mathcal{P}}\bigl{\|}Q_{\mathsf{P}}^{\pi}(k\mathbf{e})\bigr{\|}_{\rm ext}+\epsilon\inf_{c}\|k\mathbf{e}-c\mathbf{e}\|_{\rm ext}=0+0=0$ (53)
On the other hand, if $x\notin\{c\mathbf{e}:c\in\mathbb{R}\}$, we know that
$\|x\|_{\mathcal{P}}\geq\epsilon\inf_{c}\|x-c\mathbf{e}\|_{\rm ext}>0$ (54)
Thus, the kernel of $\|\cdot\|_{\mathcal{P}}$ is exactly $\{c\mathbf{e}:c\in\mathbb{R}\}$. We now show that, for any $x\in\mathbb{R}^{S}$ and $\mathsf{P}\in\mathcal{P}$,
$\bigl{\|}\mathsf{P}^{\pi}x\bigr{\|}_{\mathcal{P}}$ $=\sup_{Q\in\ Q_{\mathcal{P}}^{\pi}}\bigl{\|}Q\mathsf{P}^{\pi}x\bigr{\|}_{\rm ext}+\epsilon\inf_{c}\bigl{\|}\mathsf{P}^{\pi}x-c\mathbf{e}\bigr{\|}_{\rm ext}$
$=\sup_{Q\in\ Q_{\mathcal{P}}^{\pi}}\bigl{\|}QQ_{\mathsf{P}}^{\pi}x+QE_{\mathsf{P}}x\bigr{\|}_{\rm ext}+\epsilon\inf_{c}\bigl{\|}Q_{\mathsf{P}}^{\pi}x+E_{\mathsf{P}}x-c\mathbf{e}\bigr{\|}_{\rm ext}$
$=\sup_{Q\in\ Q_{\mathcal{P}}^{\pi}}\bigl{\|}QQ_{\mathsf{P}}^{\pi}x\|_{\rm ext}+\epsilon\inf_{c}\bigl{\|}Q_{\mathsf{P}}^{\pi}x-c\mathbf{e}\bigr{\|}_{\rm ext}$
$\leq\alpha\sup_{Q\in\ Q_{\mathcal{P}}^{\pi}}\bigl{\|}Qx\|_{\rm ext}+\epsilon\|Q^{\pi}x\|_{\rm ext}$
$=(\alpha+\epsilon)\|Q^{\pi}x\|_{\rm ext}$
$\leq(\alpha+\epsilon)\|x\|_{\mathcal{P}}.$ (55)
Since $\alpha\in(0,1)$ and $\epsilon\in(0,1-\alpha)$. Thus, let $\gamma=\alpha+\epsilon$ then $\gamma\in(0,1)$. Substituting the above result back to (C.2), we obtain
$\|\mathbf{T}_{g}(V_{1})-\mathbf{T}_{g}(V_{2})\|_{\mathcal{P}}\leq\|\pi(a|s)\tilde{p}_{(V_{1},V_{2})}\bigl{[}V_{1}(s^{\prime})-V_{2}(s^{\prime})\bigr{]}\|_{\mathcal{P}}\leq\gamma\|V_{1}-V_{2}\|_{\mathcal{P}}$ (56)
Appendix D Biased Stochastic Approximation Convergence Rate
In Section 3, we established that the robust Bellman operator is a contraction under the semi-norm $\|\cdot\|_{\mathcal{P}}$, ensuring that policy evaluation can be analyzed within a well-posed stochastic approximation framework. However, conventional stochastic approximation methods typically assume unbiased noise, where variance diminishes over time without introducing systematic drift. In contrast, the noise in robust policy evaluation under TV and Wasserstein distance uncertainty sets exhibits a small but persistent bias, arising from the estimators of the support functions $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ (discussed in Section 4). This bias, if not properly addressed, can lead to uncontrolled error accumulation, affecting the reliability of policy evaluation. To address this challenge, this section introduces a novel analysis of biased stochastic approximation, leveraging properties of dual norms to ensure that the bias remains controlled and does not significantly impact the convergence rate. Our results extend prior work on unbiased settings and provide the first explicit finite-time guarantees, which is further used to establish the sample complexity of policy evaluation in robust average reward RL. Specifically, we analyze the iteration complexity for solving the fixed equivalent class equation $H(x^{*})-x^{*}\in\overline{E}$ where $\overline{E}\coloneqq\{c\mathbf{e}:c\in\mathbb{R}\}$ with $\mathbf{e}$ being the all-ones vector. The stochastic approximation iteration being used is as follows:
with $\eta_{t}>0$ being the step-size sequence. We assume that there exist $\gamma\in(0,1)$ such that
$\|H(x)-H(y)\|_{\mathcal{P}}\leq\gamma\,\|x-y\|_{\mathcal{P}},\quad\forall x,y$ (58)
We also assume that the noise terms $\omega^{t}$ are i.i.d. and have bounded bias and variance
$\mathbb{E}[\,\|w^{t}\|_{\mathcal{P}}^{2}|\mathcal{F}^{t}]\leq A+B\,\|x^{t}-x^{*}\|_{\mathcal{P}}^{2}\quad\text{and}\quad\bigl{\|}\mathbb{E}[\,w^{t}|\mathcal{F}^{t}]\bigr{\|}_{\mathcal{P}}\leq\varepsilon_{\mathrm{bias}}$ (59)
Theorem 8.
$\mathbb{E}\Bigl{[}\|x^{T}-x^{*}\|^{2}_{\mathcal{P}}\Bigr{]}\leq\operatorname*{\mathcal{O}}\left(\frac{1}{T^{2}}\right)\|x^{0}-x^{*}\|^{2}_{\mathcal{P}}+\operatorname*{\mathcal{O}}\left(\frac{A}{(1-\gamma)^{2}T}\right)+\operatorname*{\mathcal{O}}\left(\frac{x_{\mathrm{sup}}\varepsilon_{\text{bias}}\log T}{1-\gamma}\right)$ (60)
where $x_{\mathrm{sup}}\coloneqq\sup_{x}\|x\|_{\mathcal{P}}$ is the upper bound of the $\|\cdot\|_{\mathcal{P}}$ semi-norm for all $x^{t}$.
Theorem 8 adapts the analysis of (Zhang et al., 2021) and extends it to a biased i.i.d. noise setting. To manage the bias terms, we leverage properties of dual norms to bound the inner product between the error term and the gradient, ensuring that the bias influence remains logarithmic in $T$ rather than growing unbounded, while also carefully structuring the stepsize decay to mitigate long-term accumulation. This results in an extra $\varepsilon_{\mathrm{bias}}$ term with logarithmic dependence of the total iteration $T$.
@@ Appendix D Biased Stochastic Approximation Convergence Rate
We perform analysis of the biased‐noise extension to the semi-norm stochastic approximation (SA) problem by constructing a smooth convex semi-Lyapunov function for forming the negative drift Zhang et al. (2021); Chen et al. (2025) and using properties in dual norms for managing the bias.
D.1 Proof of Theorem 8
D.1.1 Setup and Notation.
@@ D.1 Proof of Theorem 8
In this section, we override the notation of the semi-norm $\|\cdot\|_{\mathcal{P}}$ by re-writing it as the norm $\|\cdot\|_{\mathcal{N}}$ (defined in (64)) to the equivalence class of constant vectors. For any norm $\|\cdot\|_{c}$ and equivalent class $\delta_{\overline{E}}$, define the indicator function $\delta_{\overline{E}}$ as
@@
then by Zhang et al. (2021), the semi-norm induced by norm $\|\cdot\|_{c}$ and equivalent class $\overline{E}$ is the infimal convolution of $\|\cdot\|_{c}$ and the indicator function $\delta_{\overline{E}}$ can be defined as follows
@@
Where $\ast_{\inf}$ denotes the infimal convolution operator. Throughout the remaining section, we let $\overline{E}\coloneqq\{c\mathbf{e}:c\in\mathbb{R}\}$ with $\mathbf{e}$ being the all-ones vector. Since $\|\cdot\|_{\mathcal{P}}$ constructed in (50) is a semi-norm with kernel being $\overline{E}$, we can construct a norm $\|\cdot\|_{\mathcal{N}}$ such that
$\|x\|_{\mathcal{N},\overline{E}}\coloneqq(\|\cdot\|_{\mathcal{N}}\ast_{\inf}\delta_{\overline{E}})(x)=\|x\|_{\mathcal{P}}$ (63)
We construct $\|\cdot\|_{\mathcal{N}}$ as follows:
$\|x\|_{\mathcal{N}}\coloneqq\sup_{\mathsf{P}\in\mathcal{P}}\bigl{\|}Q_{\mathsf{P}}^{\pi}x\bigr{\|}_{\rm ext}+\epsilon\inf_{c\in\mathbb{R}}\bigl{\|}x-c\mathbf{e}\bigr{\|}_{\rm ext}+\epsilon\|x\|_{\rm ext}$ (64)
where $Q_{\mathsf{P}}^{\pi}$ and $\epsilon$ are defined in (50).
Lemma 8.
The operator $\|\cdot\|_{\mathcal{N}}$ defined in (64) is a norm satisfying (63).
Proof.
We first verify that $\|\cdot\|_{\mathcal{N}}$ is a norm. Regarding positivity, since all terms in (64) are non-negative, $\|x\|_{\mathcal{N}}\geq 0$ for all $x\in\mathbb{R}^{S}$ and $\|\mathbf{0}\|_{\mathcal{N}}=0$. If $x\neq\bf{0}$, since $\|\cdot\|_{\rm ext}$ is a valid norm and $\epsilon>0$, we have
$\|x\|_{\mathcal{N}}\geq\sup_{\mathsf{P}\in\mathcal{P}}\bigl{\|}Q_{\mathsf{P}}^{\pi}x\bigr{\|}_{\rm ext}+\epsilon\|x\|_{\rm ext}>0.$
Regarding homogeneity, For any $\lambda\in\mathbb{R}$, we have
$\|\lambda x\|_{\mathcal{N}}$ $=\sup_{\mathsf{P}\in\mathcal{P}}\bigl{\|}Q_{\mathsf{P}}^{\pi}(\lambda x)\bigr{\|}_{\rm ext}+\epsilon\inf_{c}\bigl{\|}(\lambda x)-c\mathbf{e}\bigr{\|}_{\rm ext}+\epsilon\|\lambda x\|_{\rm ext}$
$=|\lambda|\sup_{\mathsf{P}\in\mathcal{P}}\bigl{\|}Q_{\mathsf{P}}^{\pi}x\bigr{\|}_{\rm ext}+\epsilon|\lambda|\inf_{c}\bigl{\|}x-c\mathbf{e}\bigr{\|}_{\rm ext}+\epsilon|\lambda|\|x\|_{\rm ext}$
$=|\lambda|\|x\|_{\mathcal{N}}$
Regarding triangle inequality, for any $x,y\in\mathbb{R}^{S}$, we have
$\|x+y\|_{\mathcal{N}}$ $=\sup_{\mathsf{P}\in\mathcal{P}}\bigl{\|}Q_{\mathsf{P}}^{\pi}(x+y)\bigr{\|}_{\rm ext}+\epsilon\inf_{c}\bigl{\|}(x+y)-c\mathbf{e}\bigr{\|}_{\rm ext}+\epsilon\|x+y\|_{\rm ext}$
$\leq\sup_{\mathsf{P}\in\mathcal{P}}\|Q^{\pi}_{\mathsf{P}}x\|_{\rm ext}+\sup_{\mathsf{P}\in\mathcal{P}}\|Q^{\pi}_{\mathsf{P}}y\|_{\rm ext}+\epsilon\inf_{a,b}\bigl{\|}x-a\mathbf{e}+y-b\mathbf{e}\bigr{\|}_{\rm ext}+\epsilon(\|x\|_{\rm ext}+\|y\|_{\rm ext})$
$\leq\sup_{\mathsf{P}\in\mathcal{P}}\|Q^{\pi}_{\mathsf{P}}x\|_{\rm ext}+\sup_{\mathsf{P}\in\mathcal{P}}\|Q^{\pi}_{\mathsf{P}}y\|_{\rm ext}+\epsilon\inf_{a}\bigl{\|}x-a\mathbf{e}\bigr{\|}_{\rm ext}+\epsilon\inf_{b}\bigl{\|}y-b\mathbf{e}\bigr{\|}_{\rm ext}+\epsilon\|x\|_{\rm ext}+\|y\|_{\rm ext}$
$=\|x\|_{\mathcal{N}}+\|y\|_{\mathcal{N}}.$
We now show that since $Q^{\pi}_{\mathsf{P}}\mathbf{e}=0$ for all $\mathsf{P}\in\mathcal{P}$, by the definition of infimal convolution, we have that for all $x\in\mathbb{R}^{S}$,
$(\|\cdot\|_{\mathcal{N}}\ast_{\inf}\delta_{\overline{E}})(x)$ $=\inf_{k\in\mathbb{R}}\|x-k\mathbf{e}\|_{\mathcal{N}}$
$=\inf_{k\in\mathbb{R}}\left(\sup_{\mathsf{P}\in\mathcal{P}}\bigl{\|}Q_{\mathsf{P}}^{\pi}x-kQ_{\mathsf{P}}^{\pi}\mathbf{e}\bigr{\|}_{\rm ext}+\epsilon\inf_{c\in\mathbb{R}}\bigl{\|}x-c\mathbf{e}-k\mathbf{e}\bigr{\|}_{\rm ext}+\epsilon\|x-k\mathbf{e}\|_{\rm ext}\right)$
$=\inf_{k\in\mathbb{R}}\left(\sup_{\mathsf{P}\in\mathcal{P}}\bigl{\|}Q_{\mathsf{P}}^{\pi}x\bigr{\|}_{\rm ext}+\epsilon\inf_{c\in\mathbb{R}}\bigl{\|}x-c\mathbf{e}\bigr{\|}_{\rm ext}+\epsilon\|x-k\mathbf{e}\|_{\rm ext}\right)$
$=\sup_{\mathsf{P}\in\mathcal{P}}\bigl{\|}Q_{\mathsf{P}}^{\pi}x\bigr{\|}_{\rm ext}+\epsilon\inf_{c\in\mathbb{R}}\bigl{\|}x-c\mathbf{e}\bigr{\|}_{\rm ext}+\inf_{k\in\mathbb{R}}\left(\epsilon\|x-k\mathbf{e}\|_{\rm ext}\right)$
$=\|x\|_{\mathcal{P}}$
@@
$\|H(x)-H(y)\|_{\mathcal{N},\overline{E}}\leq\gamma\|x-y\|_{\mathcal{N},\overline{E}},\quad\gamma\in(0,1),\quad\forall x,y$ (65)
@@
  • $\mathbb{E}[\,\|w^{t}\|_{\mathcal{N},\overline{E}}^{2}|\mathcal{F}^{t}]\leq A+B\|x^{t}-x^{*}\|_{\mathcal{N},\overline{E}}^{2}$ (In the robust average reward TD case, $B=0$).
  • $\bigl{\|}\mathbb{E}[w^{t}|\mathcal{F}^{t}]\bigr{\|}_{\mathcal{N},\overline{E}}\leq\varepsilon_{\mathrm{bias}}$.
@@
Note that beside the bias in the noise, the above formulation and assumptions are identical to the unbiased setups in Section B of Zhang et al. (2021). Thus, we emphasize mostly on managing the bias.
D.1.2 Semi‐Lyapunov $M_{\overline{E}}(\cdot)$ and Smoothness.
By (Zhang et al., 2021, Proposition 1–2), using the Moreau envelope function $M(x)$ in Definition 2.2 of Chen et al. (2020), we define
@@
$c_{l}M_{\overline{E}}(x)\leq\frac{1}{2}\|x\|_{\mathcal{N},\overline{E}}^{2}\leq c_{u}M_{\overline{E}}(x),$ (67)